### v1

# Improving LLM Personas via Rationalization with Psychological Scaffolds

Abstract
Language models prompted with a user description or persona can predict a user’s preferences and opinions, but existing approaches to building personas—based solely on a user’s demographic attributes and/or prior judgments—fail to capture the underlying reasoning behind said user judgments. We introduce PB&J (Psychology of Behavior and Judgments), a framework that improves LLM personas by incorporating rationales of why a user might make specific judgments. These rationales are LLM-generated, and aim to reason about a user’s behavior on the basis of their experiences, personality traits or beliefs. This is done using psychological scaffolds—structured frameworks grounded in theories such as the Big 5 Personality Traits and Primal World Beliefs—that help provide structure to the generated rationales. Experiments on public opinion and movie preference prediction tasks demonstrate that LLM personas augmented with PB&J rationales consistently outperform methods using only a user’s demographics and/or judgments. Additionally, LLM personas constructed using scaffolds describing user beliefs perform competitively with those using human-written rationales.
Improving LLM Personas via Rationalization
with Psychological Scaffolds
Brihi Joshi Xiang Ren Swabha Swayamdipta Rik Koncel-Kedziorski Tim Paek University of Southern California  Apple {brihijos}@usc.edu
1 Introduction
Figure 1: Overview of the PB&J framework: A base persona comprising user Demographics and Judgments is augmented with post-hoc rationales of various Scaffolds generated by an LLM. The updated persona is integrated into the system prompt of an LLM to align predictions with user behavior.
Recent advances in prompting large language models (LLMs) to adopt specific user personas offer straightforward yet powerful ways to simulate human behavior and decision-making. A persona is a representation of a single user’s characteristics, preferences, and behavior, designed to guide and align interactions in a personalized system D’Onofrio (2020). Personas have enabled simulation capabilities in LLMs, such as aiding survey design by simulating a wide range of human responses Argyle et al. (2023); Santurkar et al. (2023); Tjuatja et al. (2023), simulating communities to study behavior Park et al. (2022); Zhou et al. (2024); Park et al. (2024), and producing large-scale and diverse synthetic data Moon et al. (2024); Ge et al. (2024). Yet, simulated LLM personas still struggle to align well with intended user behavior Gupta et al. (2024); Liu et al. (2024).
A typical task for evaluating user-LLM persona alignment is opinion prediction. For example, consider opinion-based questions such as “How much are you interested in visiting other countries?”, or “For your job or career aspects of your life, would you say that you are where you expected to be at this point in your life?” A well-aligned LLM persona for a particular user should be able to generate the same responses to these questions as the user. Some approaches to create these personas involve fine-tuning LLMs on user conversations and interaction history Mazaré et al. (2018); Madotto et al. (2019); Li et al. (2024); however, this approach does not scale well due to the lack of adequate user-specific data and the need to fine-tune LLMs for every new user. Zero-shot prompting approaches, where a persona is specified using demographic attributes in a system prompt, avoid fine-tuning but often produce inconsistent responses and expose known biases in current LLMs Santurkar et al. (2023); Hu and Collier (2024); Cheng et al. (2023a); Gupta et al. (2024).
To address these issues, recent approaches incorporate user judgments—prior user responses and history—in the persona prompt Hwang et al. (2023); Sun et al. (2024). For example, a user’s response of “No, not right now” to a question “Are you currently focused on your professional life and career?” is a judgment that could inform their response to the question about where they expected themselves to be career-wise at this moment. While these methods provide additional context, they still fail to address a critical gap: the ability to rationalize why a user might hold a specific judgment, an essential component for understanding behavior. In the context of our example, a rationalization for the user’s answer may be that the user is prioritizing their family over their career at this time. We hypothesize that incorporating post-hoc rationalization for why a user might have made a particular judgment can improve persona prompting in LLMs, bridging the gap between surface-level personas and deeper, more nuanced simulated behavior. However, collecting these rationales is expensive and existing datasets used for persona building are not equipped with them; they only contain user judgments Kirk et al. (2024); Durmus et al. (2024), and in rare cases, user demographics Santurkar et al. (2023); Harper and Konstan. (2015).
In this paper, we test our hypothesis by introducing PB&J (Psychology of Behaviour and Judgements), a framework that uses post-hoc rationales for user judgments, generated by LLMs, to improve LLM personas. PB&J incorporates psychological theories, such as personality traits (e.g., Big 5 traits Goldberg (1993), Schwartz’s value theory Schwartz (1992)) and belief systems (e.g., Primal World Beliefs Clifton et al. (2019)) and uses these theories to guide LLM generated-rationales for given user judgments; we call these psychological scaffolds. It is important to note that these rationales may not reflect the true reasoning behind a user’s judgment, as that is still not present in user information in available datasets. However, we hypothesize that these generated rationales contain plausible markers of real behavioral reasoning due to the presence of such reasoning in the LLM’s training data Binz and Schulz (2023); Hagendorff (2023), as people regularly use everyday reasoning to rationalize behavior — a process known as folk psychology Churchland and Haldane (1988); Malle and Knobe (1997); Malle (2004).
Given a user’s demographics and prior judgments, PB&J uses LLM-generated rationales for user’s judgments to construct a richer, more comprehensive user description for persona prompting. We hypothesize that such richer descriptions help LLMs to move away from making generalizations to more individual responses while predicting user responses Gupta et al. (2024); Hwang et al. (2023). This persona is then used by LLMs to generate predictions which better align with the user’s true responses. We evaluate PB&J on two tasks that represent different aspects of personalization: opinion prediction in OpinionQA Santurkar et al. (2023) and movie preference prediction in MovieLens Harper and Konstan. (2015). Our experiments on both datasets show that LLM personas prompted with PB&J rationales are significantly better at predicting a user’s judgments than those prompted with user demographics and/or prior judgments alone. The improvements from PB&J are seen across nearly all demographic categories, and PB&J can improve personas with or without demographic information. Additionally, we present a pilot experiment where we collect human-written rationales, along with their demographics and judgments on a subset of OpinionQA; LLM-rationales nearly match the performance of human-written rationales in PB&J, showing the efficacy of generated rationales in our approach.
2 Background
We first introduce the notion of rationalization in the context of reasoning about user behavior, similar to how humans use common explanations to rationale each other’s actions Malle (2004). We hypothesize that rationales are common enough in our discourse to be prevalent in training corpora for LLMs to be able to be used in zero-shot settings. In order to guide LLMs better, we employ various psychological scaffolds, which are templates based on more structured theories of psychology, to rationalize user judgment.
2.1 Rationalization
Psychologists have long studied how people utilize everyday language to rationalize and anticipate behavioral patterns in social interactions, otherwise known as folk psychology Churchland and Haldane (1988); Malle and Knobe (1997); Malle (2009); Gordon (1986). Building on the example introduced in Section 1, suppose that we have to rationalize why a user believes that they are not focused on their professional life and career at the moment (See Figure 1). We could attribute this behavior to some belief, such as life is dull and uninteresting and chasing career progressions is a waste of time Dennett (1981). Or, we might offer life experiences that shape beliefs, such as personal life issues that are prioritized over career at the moment Malle (2004). For our purposes, we define rationalization as an attempt to understand the behavior of others using explanations based on attributed mental states.
We hypothesize that LLMs have embedded key reasoning required to rationalize user behavior—often present in narrative text (e.g., novels, biographies, social media)—and can thus generate effective rationales Binz and Schulz (2023); Hagendorff (2023). Unlike prior work, which uses rationalization as reasoning chains to explain model decisions post-hoc—either to improve model task performance Wei et al. (2023); Zelikman et al. (2022); Ramnath et al. (2024) or enhance human utility Joshi et al. (2023); Si et al. (2024); Chaleshtori et al. (2024)—we focus on rationalization as a framework for modeling and predicting human behavior.
To rationalize with an LLM, we provide a persona $\mathcal{Q}$ and user’s judgment $j^{i}$ as input and instruct the model to generate “a reasonable explanation that the person would provide for holding that judgment”. The basic rationale $r^{i}_{\mathcal{R}}$ is defined as follows:
$r^{i}_{\mathcal{R}}=LLM_{\mathcal{R}}(j^{i},\mathcal{Q})$ (1)
Here $LLM$ is an off-the-shelf pretrained language model and $\mathcal{R}$ denotes the instructions used to elicit a rationale.
2.2 Psychological scaffolds
Scaffold Name Description
No Scaffold Free-form rationales without any structured guidance which rely on the model’s internal knowledge of folk psychology.
Experiences McAdams (1993) Rationales describing the life experiences and societal norms that would motivate a user’s judgment, such as relationships, work, or community.
Big 5 Personality Traits Goldberg (1993) Rationales use personality traits such as openness, conscientiousness, extraversion, agreeableness, and neuroticism to explain judgments.
Schwartz Theory of Basic Human Values Schwartz (1992) Rationales that ground judgments in the user’s relationship to stimulation, hedonism, self-direction, universalism, and security, among others.
Primal World Beliefs Clifton et al. (2019) Rationales which use worldview — whether the world is good, safe, alive, or enticing — to explain user judgments.
Table 1: Summary of the scaffolds used in PB&J, highlighting their theoretical underpinnings and practical relevance.
While Equation 1 helps us generate rationales from LLMs, it is important to control what LLMs generate as plausible reasons for human behavior. We believe that LLMs can provide structured rationales that incorporate well-established psychological theories. We call these psychological scaffolds as they extend the LLMs’ reasoning power and reach by providing coherent, theoretical scaffolds for reasoning about observed behavior.
The specific scaffolds we study in this work are shown in Table 1. In the context of our previous example, the experiences scaffold might promote a rationale based on the bad experience of a family member who prioritized career goals over personal life McAdams (1993). Rationales generated from the personality traits scaffold may posit the user’s openness to new experiences, including bold career choices Goldberg (1993); Schwartz (1992). A rationale using the Primal World Beliefs scaffold may surmise that the user sees the world as dull and uninteresting and thus not worth the effort of seeking professional success Clifton et al. (2019). More details about the theories underlying each scaffold can be found in Appendix H. These scaffolds offer diverse yet complementary ways to structure psychological rationales, anchoring explanations in personal experiences, personality traits, and belief systems. By providing a systematic way to interpret behavior, we show that scaffolds enhance the coherence and depth of rationalization.
To generate a rationale from an LLM with a particular scaffold, we include additional scaffold information in the rationale generation instructions to arrive at scaffold-specific instructions $\psi$. Using $\mathcal{Q}$ and $j^{i}$ as above, the scaffold rationale $r^{i}_{\psi}$ is defined as follows:
$r^{i}_{\psi}=LLM_{\psi}(j^{i},\mathcal{Q})$ (2)
The specific instructions used for each scaffold are given in Appendix A.
3 LLM Personas using PB&J
Consider a task where an LLM has to adopt the persona of a given user, and perform a task (e.g. answer opinion-based questions) as their proxy. In this section, we describe how we leverage LLM rationalization to construct LLM personas, which are used as prompts to generate personalized responses as a proxy for the user. We begin by setting up a basic persona description composed of user demographics and augmented with a set of user judgments as in previous work Hwang et al. (2023). We then present PB&J, which enriches persona descriptions by providing additional context in the form of rationales, and demonstrate how PB&J is used to generate personalized responses as a proxy for the user.
3.1 Persona Building
Base Persona Setup.
Following Hwang et al. (2023), we begin with a basic persona $\mathcal{Q}_{B}$ for an arbitrary user consisting of demographic attributes $\mathcal{D}$ and a set of $N$ seed judgments $\mathcal{J}$. Demographic attributes $\mathcal{D}$ capture the sociological characteristics and group identity traits of a user, such as age, gender, education, and race Santurkar et al. (2023). While these attributes can provide useful information on user preferences and opinions, prior work has shown that relying solely on demographics often results in stereotypical and biased responses from LLMs Hwang et al. (2023); Cheng et al. (2023b); Gupta et al. (2024). To address these limitations, Hwang et al. (2023) augment demographic attributes with a set of seed judgments $\mathcal{J}$. These seed judgments often represent user-selected answers from a held out set of questions, or a prior interaction history which aims to provide additional context to enrich the persona description Hwang et al. (2023); Do et al. (2024). As illustrated in Figure 1, the user’s response to the question “Are you currently focused on your professional life and career?” is a judgment. In other tasks, a user’s interaction history, such as the movies they watched beforehand along with their reviews, would be seed judgments.
Post-hoc Rationales for Seed Judgments.
In PB&J, we augment the above persona description $\mathcal{Q}_{B}$ with additional context in the form of post-hoc generated rationales $r^{i}$ for each seed judgment $j^{i}\in\mathcal{J}$ following Equation 1. Each post-hoc rationale is generated independently for the corresponding seed judgment and is added to the persona alongside the judgment itself. Our rationalized persona is defined as:
$\mathcal{Q}_{\mathcal{R}}=(\mathcal{D};\{(j^{i},r^{i}_{\mathcal{R}}):0\leq i\leq N\}).$ (3)
Building on Section 2.1, post-hoc rationales operationalize how $LLM_{\mathcal{R}}$ generates psychological rationales to justify why a user might hold a given judgment. Assuming that LLMs are trained on a sufficiently large corpora of human-written text—likely containing a surfeit of human rationales about behavior—they should be able to generate similar language constructs to rationalize user behavior without requiring anthropomorphism. [^1] This does not imply that LLMs actually understand user behavior; it is possible that LLMs mimic reasoning patterns found in their training data.
If we use Equation 2 to generate rationales instead, we arrive at the scaffold rationalized persona:
$\mathcal{Q}_{\mathcal{\psi}}=(\mathcal{D};\{(j^{i},r^{i}_{\mathcal{\psi}}):0\leq i\leq N\}).$ (4)
We observe that strong language models tend to follow the rationalization instructions provided in $\psi$.
3.2 Generating responses from Persona LLMs
Similar to prior work Santurkar et al. (2023); Hwang et al. (2023), persona descriptions $\mathcal{Q}$ obtained in Section 3.1 can be used as prompts to steer responses towards a given user persona. Consider a task where an individual user has provided several input-output pairs $\{x^{i},y^{i}\}$ (e.g., opinion question and answer pairs). The persona description can provide additional context to these inputs, enabling an LLM to generate predictions that align with the user’s preferences and behavior.
We predict a user’s output for an input $x^{i}$ by providing task-specific instructions $\mathcal{T}$ to an off-the-shelf language model along with the user’s persona $Q$.
$\hat{y}^{i}=LLM_{\mathcal{T}}(x^{i},\mathcal{Q}),$ (5)
To evaluate a persona building method, we vary $\mathcal{Q}$ over the base, rationalized, and scaffold rationalized personas $\mathcal{Q}_{B},\mathcal{Q}_{\mathcal{R}},\mathcal{Q}_{\psi}$. We consider aggregate prediction accuracy under this model across test users as a measure of the fidelity of the persona building method.
4 Experiments
{mdframed}
[linecolor=black!30,backgroundcolor=black!5] OpinionQA
Question: How well, if at all, do the following words or phrases describe you? Interested in visiting other countries?
Options:
A. Describes me well
B. Does not describe me well
C. Refused to answer
User-selected Answer: A. Describes me well
{mdframed}
[linecolor=black!30,backgroundcolor=red!5] MovieLens
Question: Out of 5, what would ’To Kill a Mockingbird (1962)’ be rated?
Synopsis: In small-town Alabama in 1932, Atticus Finch (Gregory Peck) is a lawyer and a widower. He has two young children, Jem and Scout. Atticus Finch is currently defending Tom Robinson, a Black man accused of raping a white woman. Meanwhile, Jem and Scout are intrigued by their neighbors, the Radleys, in particular the mysterious, seldom-seen Boo Radley.
Directors: Robert Mulligan
Cast: Gregory Peck, John Megna, Frank Overton, Rosemary Murphy, Ruth White
User-selected Answer: I would rate ’To Kill a Mockingbird (1962)’ a 2 out of 5.
Figure 2: Example Task: Shown here are example inputs and outputs for OpinionQA and MovieLens datasets respectively. Each of these instances have corresponding user selected answers.
Task and Datasets.
We conduct experiments on two datasets - OpinionQA Santurkar et al. (2023) and MovieLens Harper and Konstan. (2015). OpinionQA is a collection of American public opinion surveys conducted by PEW Research containing user-selected answers to multiple-choice questions in 15 different topics ranging from food safety to guns. The dataset also contains multiple demographic attributes for every user. Our task is to predict a user’s answer for a question, given their persona description. For OpinionQA, we use a subset of 750 users along with 10 test questions for each user. MovieLens is a movie rating dataset containing timestamped movie ratings corresponding to individual users. While the demographic information in MovieLens is not as rich as OpinionQA, we use an older subset of the dataset that records the gender, age, occupation and location of each user. Our task is to predict a user’s rating, between 1-5, for a movie, given their persona description. For MovieLens, we use a subset of 100 users, with 10 test movies for each user. Further details and splits of the dataset are presented in Appendix B. Example inputs-outputs of each of these datasets are shown in Figure 2.
Judgment: I am somewhat dissatisfied with my current job or career.
No Scaffold
At my age, I’ve had many experiences in my working life and feel that I could have achieved more. While I continue to work because I believe in keeping busy and maintaining a sense of purpose, my current job doesn’t give me the satisfaction and fulfillment I’d hoped for…
Experiences
As a man in my senior years, I’ve spent decades working. I began in manual labor, but found a calling in community work and completed an associate’s degree in social work. While this transition has been rewarding, I’ve grown somewhat dissatisfied with my job mainly because of ever-increasing paperwork and bureaucracy. They are detracting from the direct, meaningful interactions that I so enjoy and that rekindled my career passion in the first place. As I age, I’ve found that what I value most in work is not status or power…
Big 5 Personality Traits
This dissatisfaction may stem from extroversion and my desire for more social interaction that my current job may not provide. My conscientiousness also comes into play as I am a productive and organized individual who might feel stifled if these traits are not fully utilized in my career. At the same time, my neuroticism may cause me to overanalyze situations and magnify any negative aspects of my job…
Schwartz Theory of Basic Human Values
This opinion reflects my value of Achievement and Self-Direction. I value being successful and competent, and currently, my job is not allowing me to feel that sense of accomplishment. Additionally, I value independent actions and the ability to choose my own path, thus being dissatisfied with my current job also relates back to my sense of autonomy and need for personal fulfilment…
Primal World Beliefs
Despite being on the older side, I still have an enticement for experiencing new things and learning more. Sticking to the same old job with monotonous routines doesn’t appeal to me because it does not satisfy my curiosity or hunger for novelty. Additionally, my safe world belief also comes into play. I believe in stability and minimal threats but my current job doesn’t provide that comfort and ease that I expect at this point in my life. The constant worries and pressures create a sense of dissatisfaction. Lastly, I also believe in the alive world stance. I want a sense of purpose in what I do. I believe there is a reason for my dissatisfaction, and it’s perhaps because this isn’t the job I was meant to stay in for the rest of my life…
Table 2: Snippets of rationales generated by GPT-4 for a career satisfaction judgment according to the scaffolds used in PB&J. Full examples are shown in Appendix G, along with user demographics
Seed Judgments.
Our persona description takes as input a set of seed judgments, for which we generate rationales. For OpinionQA, these seed judgments are a held-out set of questions and a users’ set of corresponding answers to those questions. The questions and answers are then converted into declarative forms (e.g., If a user responds No to “Are you currently focused on your professional life and career?”, the declarative form would be I’m not very focused on my professional life right now). For MovieLens, we use movie ratings provided by the user with earlier timestamps for predicting ratings by the same user with later timestamps. Each judgment consists of a movie, its rating, and a short description of the movie in terms of its plot synopsis, actors, and directors. For our main experiments, a fixed set of 8 seed judgments are provided in the persona description.
Evaluation.
We use dataset opinion questions and movies that are not a part of the seed judgments to evaluate the personas. Since our personas are customized to each real user, we calculate accuracies and standard deviation for both OpinionQA and MovieLens, macro-averaged for each user.
Models.
We use two models of varying sizes — GPT-4  [^2] Last accessed on 2 December 2024, GPT-4 points to gpt-4-0613.OpenAI et al. (2024) and Mistral 0.2 Instruct 7B Jiang et al. (2023). All results displayed are based on prompting these models, without any fine-tuning.
Baselines.
Following prior work, we experiment with different variants of building persona descriptions as baselines, which are mostly ablations of PB&J. Santurkar et al. (2023) propose two methods – one where LLMs are prompted without any user information (No Persona) and with demographics of the user (Only Demographics). Hwang et al. (2023) add judgments to the persona descriptions without (Only Judgments) and with demographics (Demographics + Judgments). For the latter setting, we include results using Chain-of-thought reasoning Wei et al. (2023) as $\textsc{Demographics + Judgments}_{CoT}$.
Prompting Approach OpinionQA MovieLens
GPT-4 Mistral 7B GPT-4 Mistral 7B
No Persona 20.57 $\pm$ 15.15 32.08 $\pm$ 16.83 21.89 $\pm$ 13.01 06.30 $\pm$ 9.34
Only Demographics 45.69 $\pm$ 17.46 32.97 $\pm$ 16.46 35.30 $\pm$ 18.68 22.90 $\pm$ 14.65
Only Judgments 33.63 $\pm$ 16.76 38.40 $\pm$ 16.25 32.50 $\pm$ 19.10 15.30 $\pm$ 17.97
Demographics + Judgments 49.63 $\pm$ 16.84 42.17 $\pm$ 16.70 34.80 $\pm$ 17.52 21.20 $\pm$ 17.65
$\textsc{Demographics + Judgments}_{CoT}$ 49.17 $\pm$ 17.10 31.67 $\pm$ 15.59 30.20 $\pm$ 18.11 24.20 $\pm$ 18.30
[…]
Table 3: Improved Persona Alignment with PB&J: Shown here are accuracy scores and standard deviations macro-averaged across users, for PB&J with different psychological scaffolds (No Scaffold, Experiences, Big 5 Personality Traits, Schwartz Theory of Basic Human Values, and Primal World Beliefs). PB&J consistently outperforms baselines, demonstrating the effectiveness of scaffolded rationales in improving persona alignment. Best performing method is bolded and best performing baseline is underlined. * represents results that are significantly better ($p<0.05$) than the best baseline. Full significance results are in Appendix E.
Existing persona LLM baselines are (almost) equivalent.
Table 3 presents macro-averaged accuracies for various baseline methods across models and datasets. Without any persona information (No Persona condition), both GPT-4 and Mistral 7B perform significantly worse, highlighting the necessity of incorporating some form of user context. The performance drop is particularly stark for GPT-4, likely due to its safety guardrails, which often lead to abstentions from expressing opinions or preferences Chen et al. (2023). This is evidenced by frequent responses such as, "This is a subjective question and the answer will vary" or "As an AI, I don’t have personal experiences", which contribute to its lower accuracy. [^3] This behavior does not occur for all questions but has a substantial impact on overall performance. Among other baselines, demographic attributes (Only Demographics) generally provide better performance than judgments alone (Only Judgments), especially for GPT-4. The best-performing baseline varies: While Demographics + Judgments is the strongest baseline on OpinionQA, its advantage does not generalize to MovieLens, where Only Demographics performs comparably or better. This suggests that the utility of judgments depends on the nature of the task—while they offer some alignment in opinion-based questions, they are less effective in preference-based settings like MovieLens.
PB&J provides significant improvements in aligning personas.
Across both datasets (Table 3), PB&J consistently outperforms baselines, demonstrating the effectiveness of rationales. We show an example generation of rationales in Table 2. On OpinionQA, rationale scaffolds yield 9.7%–10.8% improvements over Demographics + Judgments, with Primal World Beliefs achieving the highest accuracy, closely followed by Experiences. Even unstructured rationales (No Scaffold) surpass Demographics + Judgments, reinforcing the value of rationales in persona alignment. Trends hold across models, with GPT-4 maintaining a stronger baseline.
In MovieLens, psychological scaffolds are even more critical. While GPT-4 performs better overall, Mistral 7B benefits substantially from Schwartz Theory of Basic Human Values and Primal World Beliefs, which provide the largest gains. Unlike in OpinionQA, No Scaffold fails to improve over Demographics + Judgments, indicating that psychological scaffolds may be particularly useful for modeling subjective preferences. Notably, Primal World Beliefs consistently ranks among the top scaffolds across datasets, suggesting that belief-driven rationalization is a powerful tool for persona modeling. However, Schwartz Theory of Basic Human Values excels in MovieLens, highlighting that different tasks may benefit from different psychological frameworks.
5 Using PB&J with human-written rationales
Approach Accuracy
No Persona 08.61 $\pm$ 05.20
Only Demographics 24.28 $\pm$ 15.81
Only Judgments 21.47 $\pm$ 08.50
Demographics + Judgments 39.42 $\pm$ 11.43
[…]
Table 4: Incorporating human-written rationales in PB&J: Human-written rationales for OpinionQA judgments consistently perform baselines and all LLM-generated rationales except Primal World Beliefs.
Table 3 demonstrates that LLM-generated rationales effectively improve persona alignment in opinion prediction and movie preference tasks. However, we hypothesize that the observed improvements stem not from the LLM-generated rationales themselves, but from the additional context they provide, enabling more accurate generalization for each user. To test this, we conducted a pilot study using a subset of 30 questions from OpinionQA. We recruited 100 users, collected their demographic information and responses, and asked them to provide rationales for their first nine answers. These rationales were open-ended, allowing participants to explain their opinions based on personal experiences, beliefs, or personality traits without any imposed structure. The remaining 21 questions were used for evaluation. More details about this pilot are discussed in Section B.3
Table 4 reports results on this subset comparing baseline methods, LLM-generated rationales, and Human Written rationales using GPT-4. Human-written rationales outperform all baselines and PB&J variants, with the exception of Primal World Beliefs, where the difference is not statistically significant. This highlights the effectiveness of human-generated explanations in improving persona alignment, even with limited judgments. Additionally, the comparable performance of Primal World Beliefs rationales to human-written ones suggests that carefully designed scaffolds can help LLM-generated rationales approximate human rationale quality. These findings emphasize the role of rationalization in improving persona LLMs, whether through human input or well-constructed LLM scaffolds.
6 Discussion
Figure 3: Performance as a function of the number of user judgments: PB&J outperforms baselines across all settings, providing substantial gains even with minimal judgments. All results use GPT-4.
LLM-generated rationales help with a limited budget of judgments.
Unlike demographics, user judgments take time to collect, making it crucial to assess persona alignment with limited data (Section 4). As shown in Figure 3, PB&J with Primal World Beliefs outperforms both baselines (Only Judgments and Demographics + Judgments) across all judgment counts. Even with just two judgments, PB&J surpasses baselines far more, highlighting the value of LLM-generated rationales in low-data settings. As the number of judgments increases, the performance of all methods improves. However, the rate of improvement for all diminishes after 8 judgments. This indicates that while all methods benefit from additional user judgments, PB&J maximizes its potential gains earlier due to the contextual richness provided by LLM-generated rationales.
PB&J improves performance across demographics.
Figure 4: PB&J’s improvements over Demographics + Judgments across education, race, income, and gender: Subgroups marked with * indicate significant improvements (p < 0.05). All results use GPT-4.
A robust personalization method should improve performance across diverse user groups rather than relying on gains from specific demographics. To assess this, we compare PB&J (with Primal World Beliefs scaffolds) to Demographics + Judgments across education, race, income, and gender in Figure 4 (see full results in Figure 5 in the Appendix). PB&J consistently outperforms the baseline across all demographics. While gains are notable for users with postgraduate education and those identifying as Asian or Mixed Race, improvements extend across all education levels, racial groups, income brackets, and genders. By incorporating structured rationales, PB&J improves with diverse user perspectives, demonstrating broad effectiveness without demographic biases.
PB&J improves performance without demographics.
In cases where user demographics are unavailable or where the risks of model bias outweigh the utility of demographic information, PB&J can improve performance of personalization from judgments alone. Rationalizing Only Judgments with Primal World Beliefs results in an 11.64% absolute improvement on OpinionQA and 1.1% on MovieLens (GPT-4). While still under-performing demographic-based methods, this result indicates that deeper reasoning about user behavior alone may obviate the need for coarse and potentially biasing demographic information.
7 Related Work
Personalizing LLMs.
Recent works have used LLM personas to simulate user behavior in psycholinguistic experiments Aher et al. (2023); Karra et al. (2022b), or other social science experiments Filippas et al. (2023); Argyle et al. (2023). Specifically, the use of LLMs to simulate user responses to surveys and other value-laden scenarios, using existing user information like demographic background has been gaining increasing attention Santurkar et al. (2023); Hwang et al. (2023); Durmus et al. (2024); Chuang et al. (2024); Do et al. (2024); Sun et al. (2025); Moore et al. (2024); Dillion et al. (2023); Tjuatja et al. (2023); Balepur et al. (2025) Recently, that attention has shifted more towards synthetically augmenting persona information Moon et al. (2024); Simmons (2022) or completely synthesizing personas generated from seed human data Park et al. (2022, 2024, 2023); Ge et al. (2024).
Psychology and Personas.
Psychologists have long investigated how different sociological and psychological aspects influence personality-driven narratives of an individual McAdams (1993); Bruner (1991); Pennebaker and King (1999). Recently, researchers have focused on eliciting psychological markers for evaluating LLMs Hilliard et al. (2024); Karra et al. (2022a); Serapio-García et al. (2023). Very few works actually incorporate these principles in an LLM persona itself Moon et al. (2024); Park et al. (2024).
Reasoning and Rationalization.
Previous works have focused on generating reasoning chains or rationales by either prompting language models Wei et al. (2023); Saha et al. (2023) or learning to generate rationales by fine-tuning on such data Wiegreffe et al. (2022); Ramnath et al. (2024). Recently, there has been a surge in built-in reasoning capabilities in language models via extensive ‘learning-to-reason’ steps Zelikman et al. (2022); DeepSeek-AI et al. (2025). It is important to note that these prior works focus on reasoning about the data, by primarily asking - Why did the model produce a certain output? Instead, PB&J can be seen as reasoning about people rather than the data, and scaffolds are a way to instruct the model how to reason.
8 Conclusion
This work introduces PB&J, a framework that improves LLM personas by incorporating psychologically-inspired rationales to explain user judgments. By leveraging structured psychological scaffolds, PB&J improves LLM persona performance across diverse opinion prediction and preference modeling tasks and also at par with human-written rationales. Additionally, PB&J remains effective even with limited user judgments, highlighting its potential for real-world personalization applications, with limited user history.
Limitations
While PB&J improves personas through LLM-generated rationales, it relies solely on zero-shot prompting for both rationale generation and downstream predictions. While this allows for flexibility and adaptability across users, it may not fully capture the complexity and depth of individual reasoning. Fine-tuning LLMs on human-written rationales could further improve personalization by enabling models to learn user-specific patterns rather than relying solely on generated rationales. Additionally, while these rationales improve performance, we cannot validate their fidelity to actual user reasoning, as no ground-truth rationales are available. This limitation is inherent to synthetic-data based persona modeling Moon et al. (2024); Park et al. (2024), where plausible explanations generated by LLMs may align well with observed user behavior but not necessarily reflect the true underlying motivations.
Ethics Statement
Our study primarily evaluates PB&J on U.S.-based user populations, as both OpinionQA and our human pilot study consist of participants located in the United States. Our study was conducted under the guidance of an ethics review board. Additionally, the subset of users selected from MovieLens also resides in the U.S. While this ensures consistency in evaluation, it limits the generalizability of our findings to more diverse global populations. Future work should explore the effectiveness of PB&J across different cultural and linguistic contexts to ensure broader applicability. Since our datasets involve personal judgments on opinion-based questions, some generated responses may reflect viewpoints that could be offensive or controversial. While we do not directly intervene in the LLMs’ generation of rationales, it is crucial to recognize that models can inherit biases present in both training data and user-generated inputs. Finally, as with any system that models human behavior, there are concerns around user privacy. While our work does not use real user data beyond voluntary survey responses, deploying such approaches in real-world settings would require careful consideration of data collection practices, consent mechanisms, and safeguards against potential misuse.
Appendix A PB&J prompts
In Table 5 and Table 6, we provide the prompts that we use to generate scaffolded rationales and predictions from both GPT-4 and Mistral 7B.
System Message:
A person can be described as follows:
<demographic information>
The person has the following judgements: 1. <judgement_1>
2. <judgement_2>
3. <judgement_3>
User Message:
No Scaffold
For a given judgement, what would be a reasonable explanation that the person would provide for holding that judgement?
Judgement: <judgement_i>
Experiences
For a given judgement, what would be a reasonable explanation that the person would provide for holding that judgement? The explanation should contain a specific experience or personality trait (for example, fill in details of the university, or food, or any other detail, that the person would use to better explain their judgment).
Judgement: <judgement_i>
Big 5 Personality Traits
For a given judgement, what would be a brief, reasonable explanation that the person would provide for holding that judgement? The explanation should be grounded in the big five personality traits listed below:
1. Openness to experience (includes aspects such as intellectual curiosity and creative imagination)
2. Conscientiousness (organization, productiveness, responsibility)
3. Extroversion (sociability, assertiveness; its opposite is Introversion)
4. Agreeableness (compassion, respectfulness, trust in others)
5. Neuroticism (tendencies toward anxiety and depression)
Judgement: <judgement_i>
Schwartz Theory of Basic Human Values
For a given judgement, what would be a brief, reasonable explanation that the person would provide for holding that judgement? The explanation should be grounded in the Shwartz Theory of Basic Human Values listed below:
1. Power: Refers to the pursuit of social status, dominance, and control over people and resources.
2. Achievement: Personal pursuit of success, demonstrating competence according to social standards.
3. Hedonism: Pursuit of pleasure, enjoyment, and sensory and emotional gratification.
4. Stimulation: Seeks novelty and challenge in life, valuing excitement, variety, and adventure.
5. Self-direction: Independent thought and action — choosing, creating, and exploring.
6. Universalism: Understanding, appreciation, tolerance, and protection for the welfare of all people and nature.
7. Benevolence: Preserving and enhancing the welfare of those with whom one is in frequent personal contact (the ‘in-group’).
8. Tradition: Respect, commitment, and acceptance of the customs and ideas that traditional culture or religion provide the self.
9. Conformity: Restraint of actions, inclinations, and impulses likely to upset or harm others and violate social expectations or norms.
10. Security: Safety, harmony, and stability of society, relationships, and the self.
Judgement: <judgement_i>
Primal World Beliefs
For a given judgement, what would be a brief, reasonable explanation that the person would provide for holding that judgement? The explanation should be grounded in the three primal world beliefs listed below:
1. Safe World Belief: Those low on Safe see a Hobbesian world defined by misery, decay, scarcity, brutality, and dangers of all sorts. Those high on Safe see a world of cooperation, comfort, stability, and few threats.
2. Enticing World Belief: Those low on Enticing inhabit dull and ugly worlds where exploration offers low return on investment. Those high on Enticing inhabit an irresistibly fascinating reality.
3. Alive World Belief: Those low on Alive inhabit inanimate, mechanical worlds without awareness or intent. Those high on Alive sense that everything happens for a purpose and are thus sensitive to those purposes.
Judgement: <judgement_i>
Table 5: Prompts used to generate rationales for judgments: We use a common system message that includes a user’s demographic information and all prior judgments held by the user. The user message then includes scaffold specific instructions $\psi$ to generate rationales for a specific judgment.
System Message:
You are the following person:
<demographic information>
You have the following opinions: 1. <judgement_1>+<rationale_1>
2. <judgement_2>+<rationale_2>
3. <judgement_3>+<rationale_3>
User Message:
Based on your demographic and opinion information above, which answer would you select for the above question?
Question: <question>
Answer choices: <choice>
Table 6: Prompts used to predict answers, given a persona: We use a common system message that includes a user’s demographic information and all prior judgments, along with generated rationales. The user message then includes the exact question and answer choices.
Appendix B Dataset Details
B.1 OpinionQA
OpinionQA Santurkar et al. (2023) contains fifteen topics, with multiple questions in each topic. Users answer several questions for a topic. There is no 1:1 correspondence between users in different topics. To this end, we select 50 users per topic, resulting in 750 unique users. For each user, we separate 8 answered questions and use them as prior judgments (this number changes when we run evaluations with varying number of judgments 3). All remaining questions as used as a ‘test set’, out of which we randomly select 10 questions for evaluation. All configurations and demographic setup is used similar to Hwang et al. (2023).
B.2 MovieLens
MovieLens Harper and Konstan. (2015) dataset contains user ratings (on a scale of 1-5) for movies belonging to different genres. These ratings are time-stamped. While there are limited user demographic attributes, the dataset contains information about a user’s age, occupation, location (zipcode) and gender. We convert the zipcode of every participant to a string-based city/state location Esmukov and contributors (2025). For curating judgments, we order movie ratings based on their timestamp, and pick the first 8 ratings as judgments in the persona, and from the remaining ratings, sample 10 to be used for evaluation. The judgments include the movie name and a user’s rating. We augment the judgments with the movie’s synopsis and key cast and crew using the IMDb API Pasotti and contributors (2025).
B.3 Human Pilot Dataset
For the human pilot experiment, we recruit $100$ participants from a third-party user study platform called User Research International. We selected a subset of 30 questions from OpinionQA belonging to 3 different topics: food, economics and America in 2050. Participants who consent to the study are requested to answer all $30$ questions, but can refuse any question they want. For $9$ questions, we also ask participants to provide a free-text rationale justifying their selection.
Appendix C Model Details
Config GPT-4 Mistral 7B
model GPT-4-0613 Mistral 7B 0.2 Instruct
Number of parameters: Unknown Number of parameters: 7 billion
Rationale Generation
new_tokens 256 256
temperature 1 1
seed 6 6
GPU N/A, openai api call 3 A100
Inferring time 2 hours 1 hour
Answer Generation
new_tokens 10 275
temperature 0 0
seed 6 6
GPU N/A, openai api call 3 A100
Inferring time 2 hours 1 hour
Table 7: Model Configurations for Rationale and Answer Generation
We provide model configurations for both GPT-4 and Mistral 7B for rationale generation and answer generation steps in Table 7. GPT-4 is under proprietary license, and Mistral 7B is subject to the Apache 2.0 license.
Appendix D Additional Discussion
D.1 Does Point-of-View (PoV) matter while building LLM personas?
Approach Rationale PoV Answer PoV Accuracy
Demographics + Judgments - first 32.05
Demographics + Judgments - third 49.63
[…]
Table 8: Ablations with different Point-of-Views (PoVs) in PB&J: We experiment with different PoVs to prompt answers and rationales, for the best forming baseline, and variant of PB&J. All experiments on GPT-4.
Our analysis examines how the point of view (PoV) in rationale and answer generation affects persona alignment in PB&J. The baseline model benefits from using a third-person PoV for answers, suggesting that distancing the model from a subjective stance improves alignment. In contrast, PB&J performs best when both the rationale and answer are generated in first-person, reinforcing the effectiveness of maintaining a consistent, personalized perspective. Performance declines when either the rationale or answer shifts to third-person, indicating that while baselines may benefit from objective framing, first-person perspectives enhance persona consistency when paired with rationale-augmented personas. Therefore, all baselines depicted in Table 3 are prompted in third-person and all PB&J variants are prompted in first-person, for both the rationale and answer prompts.
Appendix E Significance Tests
Dataset Model Best Baseline Approach User-wise Significance Question-wise Significance
OpinionQA GPT-4 Demographics + Judgments No Scaffold statistic=-8.9082, $p$=1.94e-18 statistic=-9.0335, $p$=1.04e-19
Experiences statistic=-9.8689, $p$=5.57e-22 statistic=-9.8824, $p$=3.42e-23
Big 5 Personality Traits statistic=-9.1648, $p$=2.34e-19 statistic=-9.0564, $p$=8.47e-20
Schwartz Theory of Basic Human Values statistic=-9.0304, $p$=7.14e-19 statistic=-8.8108, $p$=7.62e-19
Primal World Beliefs statistic=-9.9742, $p$=2.19e-22 statistic=-9.8010, $p$=7.61e-23
Mistral 7B Demographics + Judgments No Scaffold statistic=-6.1466, $p$=6.43e-10 statistic=-6.1429, $p$=4.26e-10
Experiences statistic=-6.8153, $p$=9.68e-12 statistic=-7.2184, $p$=2.89e-13
Big 5 Personality Traits statistic=-4.4573, $p$=4.78e-06 statistic=-4.6148, $p$=2.00e-06
Schwartz Theory of Basic Human Values statistic=-3.5916, $p$=1.75e-04 statistic=-3.7695, $p$=8.24e-05
Primal World Beliefs statistic=-3.9861, $p$=3.69e-05 statistic=-4.1223, $p$=1.90e-05
MovieLens GPT-4 Only Demographics Big 5 Personality Traits statistic=-0.2627, $p$=0.397 statistic=-0.3260, $p$=0.372
Schwartz Theory of Basic Human Values statistic=-2.6476, $p$=0.00472 statistic=-3.1586, $p$=0.00082
Primal World Beliefs statistic=-1.7573, $p$=0.04098 statistic=-1.8790, $p$=0.03027
Mistral 7B Only Demographics Big 5 Personality Traits statistic=-3.0672, $p$=0.00139 statistic=-3.5543, $p$=0.00020
Schwartz Theory of Basic Human Values statistic=-1.8034, $p$=0.03718 statistic=-2.1552, $p$=0.01569
Primal World Beliefs statistic=-3.8985, $p$=8.81e-05 statistic=-4.1427, $p$=1.86e-05
Human Pilot GPT-4 Demographics + Judgments No Scaffold statistic=-4.9186, $p$=1.74e-06 statistic=-5.5086, $p$=2.03e-08
Experiences statistic=-2.4163, $p$=0.00876 statistic=-2.9501, $p$=0.00161
Big 5 Personality Traits statistic=-3.0124, $p$=0.00165 statistic=-3.5183, $p$=0.00022
Schwartz Theory of Basic Human Values statistic=-3.4111, $p$=0.00047 statistic=-4.0207, $p$=3.00e-05
Primal World Beliefs statistic=-7.2335, $p$=5.05e-11 statistic=-7.8278, $p$=3.90e-15
Human Pilot statistic=-8.7457, $p$=2.96e-14 statistic=-9.2996, $p$=1.71e-20
Table 9: Statistical significance of PB&J’s improvements for both GPT-4 and Mistral 7B: Results from one-tailed independent t-tests comparing PB&J variants to the best-performing baseline. We report both user-wise and question-wise significance.
We conduct statistical tests to assess whether the improvements of PB&J over the best-performing baseline are statistically significant. Table 9 presents significance test results, comparing each PB&J variant with the strongest baseline in its respective setting. We use a one-tailed independent t-test Student (1908) to evaluate the null hypothesis that PB&J does not provide a significant improvement over the baseline. To ensure robustness, we compute significance in two ways. The first approach, user-wise significance, examines whether PB&J improves performance on a per-user basis, assessing whether predictions for individual users show meaningful gains. The second, question-wise significance, evaluates improvements across all instances of a user, aggregating performance over multiple questions answered by the same user. For each comparison, we report the test statistic and p-value in Table 9.
Appendix F Analysis of LLM generated explanations
Approach
% of rationales
that have
atleast one
demographic
mentioned ($\uparrow$)
Demographic
Keyword
Density
(# of demographic
keywords /
rationale) ($\uparrow$)
Avg. diversity
of 2-grams
in rationales
per user ($\downarrow$)
Avg. diversity
of 2-grams
in rationales per
judgment ($\downarrow$)
[…]
Table 10: Analysis of LLM-generated explanations across psychological scaffolds, measuring demographic mentions, keyword density, and 2-gram diversity: Results suggest that strong persona alignment does not necessarily require high demographic reliance or linguistic variation. All analysis on rationales generated by GPT-4.
To better understand how different scaffolds influence the content and structure of LLM-generated rationales, we analyze explanations along three key dimensions: demographic reliance, lexical diversity, and structural repetition (Table 10). The percentage of explanations mentioning demographics and the density of demographic keywords capture how often rationales explicitly reference user identity. The average 2-gram diversity per user and per judgment reflects the linguistic variability of explanations, with lower values indicating greater diversity (i.e., less repetition). This score is also known as self-repetition Salkar et al. (2022); Shaib et al. (2024).
[…]
Figure 5: PB&J’s improvements over Demographics + Judgments across all demographics: Subgroups marked with * indicate significant improvements (p < 0.05). All results use GPT-4.
Appendix G More examples of PB&J-generated rationales
We present examples of personas constructed by PB&J. For given users with demographics, we display a representative judgment provided by the user, and present rationales generated by PB&J using different scaffolds for the same judgment. We add examples in Table 11, Table 12 and Table 13.
User Demographics:
Age: 18-29, Citizenship: Yes, Region: South, Education: Postgraduate, Income: \$75,000-\$100,000, Marital status: Married, Political ideology: Liberal, Political party: Democrat, Race: White, Religion: Nothing in particular, Frequency of religious attendance: Never, Gender: Female
Selected Judgment:
I am somewhat concerned that global climate change will harm me personally at some point in my lifetime.
PB&J-generated rationales:
No Scaffold
I believe that the effects of global climate change are increasingly evident and impactful. The rise in natural disasters, unpredictable weather, food and water shortages, even though they might seem distanced, have the potential to directly affect my lifestyle, health, and well-being. It’s a concern not just for the distant future, but for my lifetime as well, hence my concern.
Experiences
As a postgraduate scholar majoring in environmental sciences, I am constantly exposed to alarming research findings on global climate change. Personally, I love outdoor activities like hiking and camping, and it’s heartbreaking seeing the devastation in forests due to wildfires reportedly linked to climate change. This, coupled with evidence of rising sea levels and increased frequency of extreme weather events, propels my concern about how this can directly impact me and our planet.
Big 5 Personality Traits
My concern about global climate change affecting me personally stems from my high level of Conscientiousness. As someone responsible and oriented toward future planning, I understand the significant risks and consequences posed by climate change. Moreover, my Openness to experience allows me to accept and internalize scientific evidence and reports that underline the severity of the issue. My slight Neuroticism also plays a role, making me more prone to worry about potential negative outcomes, hence my concern about personal harm due to global climate change.
Schwartz Theory of Basic Human Values
This opinion is grounded in a few of Shwartz’ basic human values. Primarily, it expresses my value of Universalism, as I am aware of the interconnections and mutual dependencies in our global ecosystem and am concerned for the welfare of nature. It also links to Security, as the threat of climate change disrupts the stability of my personal life and the wider community. Lastly, the opinion slightly touches on Self-direction. My awareness and concern reflect my informed understanding and independent thought on global matters.
Primal World Beliefs
This viewpoint aligns with my low Safe World Belief, as one can view global climate change as a significant and imminent threat to stability and personal safety. Despite the advancements in technology and modern conveniences, our safety is still compromised by environmental degradation. Furthermore, my high Enticing World Belief encourages me to explore and understand more about this global issue, its impacts, and the ways to mitigate it. However, my low Alive World Belief also reinforces this concern as climate change does not seem to occur with comprehensible intent or purpose, instead contributing to an unpredictable threat to my life.
Table 11: Example 1
User Demographics:
Age: 65+,Citizenship: Yes, Region: Northeast, Education: High school graduate, Income: \$100,000 or more, Marital status: Divorced, Political ideology: Very conservative, Political party: Other, Race: White, Religion: Protestant, Frequency of religious attendance: Never, Gender: Female
Selected Judgment:
I feel no pressure from family members to marry my partner.
PB&J-generated rationales:
No Scaffold
I believe in independence and making my own decisions. Also, having been through a divorce, my family understands and respects my perspective about not wanting to remarry.
Experiences
After my divorce, my family has been exceptionally supportive of my decisions regarding my love life. With my grown children doing well and leading their own lives, they are more focused on my happiness than societal expectations. They have always wanted me to be comfortable and happy, rather than rush into any decisions. That’s why I haven’t felt any pressure from them to marry my current partner.
Big 5 Personality Traits
As a very conservative person who is also divorced, I’ve learned the importance of self-reliance and making decisions guided by my own perceptions and experiences (relating to the trait of conscientiousness). My family respects my conservative beliefs and my decision not to remarry, thereby causing no pressure for me to marry my partner. Furthermore, my trait of extroversion makes me capable of effectively communicating my decisions and viewpoints to my family, reinforcing their understanding and respect for my personal choices.
Schwartz Theory of Basic Human Values
This opinion is rooted in the values of Self-direction and Achievement. My high-income bracket has allowed me a certain degree of independence and control over my life choices, which in turn alleviates any external pressure from family to marry my partner. Moreover, being of a mature age, I have lived through the social norms and expectations often imposed on younger generations. Hence, my achievements in life so far have given me the confidence and autonomy to make my own decisions.
Primal World Beliefs
As someone with a very high Safe World Belief, I live in a cooperative and comforting environment where my decisions are respected. This harmonious setting also allows for open conversation about my personal choices, including my decision to stay unmarried to my partner. Thus, my family members do not pressure me to marry.
Table 12: Example 2
User Demographics:
Age: 30-49, Citizenship: No, Region: South, Education: College graduate/some postgrad, Income: \$50,000-\$75,000, Marital status: Married, Political ideology: Conservative, Political party: Independent, Race: Black, Religion: Protestant, Frequency of religious attendance: More than once a week, Gender: Male
Selected Judgment:
The promotion of racial and ethnic diversity in the workplace is somewhat important for companies and organizations.
PB&J-generated rationales:
No Scaffold
While I believe in the importance of individual skill and merit, I also acknowledge the role racial and ethnic diversity plays in fostering a broad range of perspectives in the workplace. This, in turn, can drive innovation, create a more inclusive environment, and better represent the diverse markets that today’s companies and organizations serve. Nevertheless, I don’t regard it as the most crucial factor for success. Hiring should always focus primarily on capability and qualifications.
Experiences
As someone who has spent a significant portion of my career working in a corporate setting, I have observed firsthand how diversity contributes to a more innovative and productive work environment. Companies with a wide range of perspectives—from people of all races, ethnicities, and backgrounds—tend to solve problems more creatively and be more adaptable in the face of change. While I believe many factors play into a business’s success, I have come to appreciate that variety in cultural background and experience certainly enriches the workplace and can provide a competitive edge.
Big 5 Personality Traits
With my level of conscientiousness, I believe that recruiting a workforce that is diverse allows companies and organizations to draw upon a broader range of insights, skills, and experiences, thereby enhancing their productivity and competitiveness. My openness to experience influences my belief that this diversity brings in new and varied perspectives which leads to innovative problem solving. However, it’s only ’somewhat’ important because I believe merit should be the primary determining factor for hiring and promotions, rather than meeting a quota system.
Schwartz Theory of Basic Human Values
This opinion aligns with my value of Universalism, as promoting racial and ethnic diversity in the workplace fosters an environment of tolerance and respect for all people, regardless of their background. Additionally, from a Power perspective, a diverse workplace may offer more opportunities for differing perspectives, which could yield more balanced and inclusive decision-making processes, thereby amplifying social status and control over resources. Furthermore, such diversity contributes to Achievement by elevating a organization’s competency and credibility in an increasingly globalized world.
Primal World Beliefs
As someone who has spent a significant portion of my career working in a corporate setting, I have observed firsthand how diversity contributes to a more innovative and productive work environment. Companies with a wide range of perspectives—from people of all races, ethnicities, and backgrounds—tend to solve problems more creatively and be more adaptable in the face of change. While I believe many factors play into a business’s success, I have come to appreciate that variety in cultural background and experience certainly enriches the workplace and can provide a competitive edge.
Table 13: Example 3
Appendix H More on Psychological Scaffolds
A rich body of research has extended, debated, and validated each of the psychological scaffolds presented earlier. Among these, Primal World Beliefs stands out as the most linguistically motivated framework, which may explain its strong performance in our experiments. As described in Clifton et al. (2019), researchers analyzed historical texts and over 80K tweets using topic modeling and concept extraction to identify statements about how people perceive the world. These statements were then categorized through expert coding, consultation with social scientists, and discussions with religious focus groups, leading to the identification of 26 fundamental “primal world beliefs". These beliefs encapsulate deep-seated assumptions individuals hold about the world, such as whether the world is inherently safe or dangerous, simple or complex, and abundant or limited. Expanding on this work, Vu et al. (2022) introduced a Latent Beliefs Model, which leverages transformer-based embeddings and a modified GPT-2 model to automatically infer latent dimensions of human beliefs from social media text. This data-driven discovery of worldviews underscores the linguistic basis of primal beliefs and their connection to naturally occurring human explanations.
Another well-established framework in social psychology is Schwartz Theory of Basic Human Values Schwartz (1992); Schwartz et al. (2012). This theory posits that human values—fundamental guiding principles—are organized along universal motivational dimensions that drive behavior. Schwartz identifies ten broad value categories, such as self-direction (independence of thought and action), benevolence (concern for others’ welfare), and power (desire for dominance or control). These values are structured in a circular model, where adjacent values are more compatible, and opposing values (e.g., security vs. stimulation) tend to be in tension. A key aspect of Schwartz’s values is their cross-cultural validation; extensive empirical studies have shown that these value dimensions hold across diverse populations, making them a robust framework for modeling user judgments in an LLM setting. Unlike primal beliefs, which describe broad worldviews, Schwartz’s values provide a structured way to infer decision-making tendencies and moral considerations, making them particularly useful for understanding user preferences in ethical or societal questions.
Similarly, the Big 5 Personality Traits (also known as the OCEAN model) Goldberg (1993); Costa and McCrae (1999); Pennebaker and Graybeal (2001); Schwartz et al. (2013) offer a comprehensive framework for describing individual differences in personality. The Big Five dimensions—Openness to Experience, Conscientiousness, Extraversion, Agreeableness, and Neuroticism—have been extensively validated through psychometric studies and natural language analysis. These traits predict a wide range of behaviors, from political preferences to purchasing decisions, and have been found to correlate with linguistic patterns in social media and personal narratives. For instance, individuals high in Openness to Experience tend to use more abstract and imaginative language, while those high in Neuroticism are more likely to express negative emotions. Given these correlations, Big Five traits serve as a useful scaffold for generating rationales that reflect personality-driven reasoning processes, such as why a highly conscientious user might favor structured decision-making or why an extraverted user might prioritize social considerations.
Each of these psychological scaffolds offers a unique perspective on human behavior: Primal World Beliefs focus on fundamental assumptions about the world, Schwartz’s values provide a structured way to model decision-making, and the Big Five capture stable personality traits that influence judgment and preference formation.
Appendix I Wasserstein Distance Metrics
Approach GPT-4 Mistral 7B
No Persona 0.95 $\pm$ 0.50 2.66 $\pm$ 0.66
Only Demographics 0.75 $\pm$ 0.42 0.97 $\pm$ 0.53
Only Judgments 0.83 $\pm$ 0.44 1.31 $\pm$ 0.62
Demographics + Judgments 0.81 $\pm$0.45 1.03 $\pm$ 0.56
$\textsc{Demographics + Judgments}_{CoT}$ 0.88 $\pm$0.39 0.97 $\pm$ 0.27
[…]
Table 14: Wasserstein Distance Metrics for MovieLens
In addition to accuracy, we also present the Wasserstein Distance (WSD) between predicted ratings and user-provided ratings for the MovieLens dataset. We present these results in Table 14. Note that lower values are better for Wasserstein Distance.